OpenStudy (anonymous):

Describe two different ways to solv 5/6=x/24 demonstrate both methods! PLZ HELP!

OpenStudy (anonymous):

 or just solve for x by multiplying both side by 24 \[5/6=x/24 ==> x = 5(24)/6 = 120/6 = 15\]